OpenStudy (anonymous):

Find the slope and y-intercept of the line 14x+4y=24

OpenStudy (anonymous):

Do you know what slope intercept form look like?

OpenStudy (anonymous):

yes I do know but it just confuses me so much trying to flip things around and I honestly don't know how to do this. and Rhystic; you honestly didn't help me.

OpenStudy (rhystic):

hmm. so you understand y = mx + b correct?

OpenStudy (anonymous):

Crosses the y-axis @ 6

OpenStudy (anonymous):

Yes I do.

OpenStudy (anonymous):

how about graph the line using this graphing calculator Desmos.com and then using two points from the line figure out the slope =)

OpenStudy (rhystic):

basically solve for y.... 14x + 4y = 24 -> 4y = 24 - 14x -> y = 24/4 - 14/4x to put it completely in y = mx + b form -> y = -(14/4)x + 6

OpenStudy (anonymous):

but in order to graph it I would need it in y=mx+b form

OpenStudy (anonymous):



OpenStudy (anonymous):

nope all you do is type that equation ^^^^^ in the calculator and it will graph it =)

OpenStudy (rhystic):

youre simply subtracting 14x from both sides and then dividing 4 on both sides to get y isolated on one side of the equation yielding a y = mx +b form.

OpenStudy (anonymous):



OpenStudy (rhystic):

defame look above a couple posts and see if you follow what I did.

iYuko (iyuko):

This might help @DefameMe http://openstudy.com/updates/5533407ce4b021d09bb190df